The volume of the box is represented by x squared plus five x plus six, times x plus five. Find the polynomial that represents the area of the bottom of the box? Height is given by x plus 2.
Answer by Earlsdon(6294)

Volume, V = (x^2 + 5x + 6)(x + 5) Factor the trinomial.
V = (x + 2)(x + 3)(x + 5)
V = (H )(L)(W) Since the height of the box is given as: (x + 2), then the length and width of the base must be: (x + 5) and (x + 3)
Area of the base = (L)(W) = (x + 5)(x + 3) = x^2 + 8x + 15
So, Area of base = x^2 + 8x + 15